A “pool shark” takes aim at the eight ball in a game of billiards. Her shot gives the cue ball a velocity of 2.5 m/s [forward]. The cue ball strikes the eight ball and continues with a velocity of 0.25 m/s [forward]. If the mass of both the cue ball and the eight ball equals 400 g, calculate the velocity of the eight ball after collision.
Answer #1

Let us consider the forward direction that is the initial velocity direction of the cue ball as positive.

Mass of the cue ball = m1 = 400 g = 0.4 kg

Mass of the eight ball = m2 = 400 g = 0.4 kg

Initial velocity of the cue ball = V1 = 2.5 m/s (Positive as it is forward)

Initial velocity of the eight ball = V2 = 0 m/s

Velocity of the cue ball after the collision = V3 = 0.25 m/s (Positive as it is forward)

Velocity of the eight ball after the collision = V4

By conservation of linear momentum,

miVi + m2V2 = miV3 + m2V4

(0.4) 2.5) + (0.4) 0) = (0.4)(0.25) + (0.4)V4

V4 = 2.25 m/s

The velocity is positive therefore the direction of the velocity is forward.

Velocity of the eight ball after the collision = 2.25 m/s forward